OpenStudy (anonymous):

The pool at a park is circular. You want to find the equation of the circle that is the boundary of the pool. Find the equation if the area of the pool is 400 square feet and (0, 0) represents the center of the pool. a. x2 + y2 = 20 b. x2 + y2 = 400 c. x2 + y2 = 400 pi d. x2 + y2 = 400/p

OpenStudy (anonymous):

can anyone help me?

jimthompson5910 (jim_thompson5910):

use the equation (x-h)^2+(y-k)^2 = r^2 where (h,k) is the center and r is the radius In this case, the center is (h,k) = (0,0). So h = 0 and k = 0 The area of the pool is 400 square ft, so the radius squared is A = pi*r^2 400 = pi*r^2 400/pi = r^2 r^2 = 400/pi Plug all this info into the first equation given to get (x-h)^2+(y-k)^2 = r^2 (x-0)^2+(y-0)^2 = 400/pi x^2 + y^2 = 400/pi So the answer is choice D